# Principal stratification with recurrent events truncated by a terminal event: A nested Bayesian nonparametric approach

### v2

@@ 3 Bayesian Nonparametric Causal Inference
Under Assumptions 14, (1) and (2) are nonparametrically identified up to distribution of the frailty $\boldsymbol{\gamma}$ as follows:
$\begin{split}\mu^{z}\quantity(t;z)&=\frac{\int_{\mathcal{X}}\int_{0}^{\infty}\kappa_{t,r}(z,x,\boldsymbol{\gamma})\eta_{r}(1,x,\boldsymbol{\gamma})\eta_{r}(0,x,\boldsymbol{\gamma})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{X}(x)d\boldsymbol{\gamma}dx}{\int_{\mathcal{X}}\int_{0}^{\infty}\eta_{r}(1,x,\boldsymbol{\gamma})\eta_{r}(0,x,\boldsymbol{\gamma})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma}^{\prime})f_{X}(x^{\prime})d\boldsymbol{\gamma}^{\prime}dx^{\prime}},\\ \tau^{z}(t;r)&=\frac{\int_{\mathcal{X}}\int_{0}^{\infty}\zeta_{t,r}(z,x,\boldsymbol{\gamma})\eta_{r}(1,x,\boldsymbol{\gamma})\eta_{r}(0,x,\boldsymbol{\gamma})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{X}(x)d\boldsymbol{\gamma}dx}{\int_{\mathcal{X}}\int_{0}^{\infty}\eta_{r}(1,x,\boldsymbol{\gamma})\eta_{r}(0,x,\boldsymbol{\gamma})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma}^{\prime})f_{X}(x^{\prime})d\boldsymbol{\gamma}^{\prime}dx^{\prime}},\end{split}$ (5)
where $\kappa_{t,r}(z,x,\boldsymbol{\gamma})=\mathbb{E}[N_{i}(t)\mid Z_{i}=z,T_{iN_{i}(r)}<D_{i},\mathbf{X}_{i}=\mathbf{x},\boldsymbol{\gamma}]$, $\zeta_{t,r}(z,x,\boldsymbol{\gamma})=\mathbb{E}[T_{iN_{i}(t)}\mid Z_{i}=z,T_{iN_{i}(r)}<D_{i},\mathbf{X}_{i}=\mathbf{x},\boldsymbol{\gamma}]$, and $\eta_{r}(z,x,\boldsymbol{\gamma})=\mathrm{Pr}(T_{iN_{i}(r)}<D_{i}\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x},\boldsymbol{\gamma})$.
The proof is provided in Web Appendix C. Theorem 1 motivates a g-computation strategy for estimating the target estimand based on observed data. This result result clarifies that the correlation parameter $\rho$ is not identifiable, which in turn suggests a sensitivity framework for posterior inference. We detail the specification of the working models in Section 3.1.
3.1 BNP for principal stratificaion with recurrent events truncated by a terminal event
@@ 3.1.1 Enriched dependent Dirichlet process
A popular BNP approach for recurrent processes with a terminal event is jointly modeling the gap times of recurrent and terminal events using a Dirichlet process (DP) mixture model. While this is a simple approach adopted in the literature for the joint analysis of recurrent events with a terminal event (Paulon2020; Xu2022; Tian2024), a potential pitfall is that the latent random partition induced by the DP will be overwhelmingly determined by the recurrent events rather than the terminal event as the number of recurrences grows, leading to unreliable within-cluster predictions of the survival time, with higher posterior variance due to unnecessarily small sample sizes. Additionally, existing approaches use the joint DP primarily to model a frailty term, but it is somewhat counterintuitive that the random partition, which yields the subject-specific frailty, is driven largely by numerous recurrent events rather than the subject-specific survival event. More technical discussions on this partitioning issue are provided in Web Appendix B.
@@
We address these issues by introducing a nested structure into the DP prior tailored for the analysis of recurrent events with semi-competing risks. The key idea is to introduce a nested partition for the unknown random joint probability measure $\mathbf{P}$. In particular, we consider the random marginal $\mathbf{P}_{\phi}$ and the random conditional $\mathbf{P}_{\theta\mid\phi}$ to obtain the desired clustering structure. Then, the nested prior is defined as $\mathbf{P}_{\phi}\sim\mathrm{DP}(\alpha_{\phi},P_{0\phi})$, $\mathbf{P}_{\theta\mid\phi}\sim\mathrm{DP}(\alpha_{\theta}(\phi)$, $P_{0\theta\mid\phi}(\cdot\mid\phi))$ for all $\phi\in\Phi$, where $\mathbf{P}_{\theta\mid\phi}$ for $\phi\in\Phi$ are independent of $\mathbf{P}_{\phi}$. These assumptions induce a prior for the random joint distribution $\mathbf{P}$ through the joint law of the marginal and conditionals and the mapping $(P_{\phi},P_{\theta\mid\phi})\to\int P_{\theta\mid\phi}(\cdot\mid\phi)dP_{\phi}$. Then, the prior is parameterized by the base measure $P_{0}$, defined by $P_{0}(A\times B)=\int_{A}P_{\theta\mid\phi}(B\mid\phi)dP_{\phi}$. This nested prior structure is called the Enriched Dirichlet process (EDP) prior (Wade2011).
@@
One core assumption of the EDP is that the distributions are exchangeable at both survival and recurrent event levels. However, in practice, we often have access to subject-specific covariates $\mathbf{X}_{i}$ and time-varying covariates $\mathbf{V}_{ij}$ at the $j$-th recurrence. These variables provide useful information in characterizing the distributions at both levels, and we hence further incorporate the dependence of the covariates on the EDP through the use of the dependent Dirichlet process (DDP) (Maceachern1999) to relax the exchangeability assumption. We refer to the resulting prior as the enriched dependent Dirichlet process (EDDP). The enriched dependent Dirichlet process mixture (EDDPM) uses the EDDP as a prior for the mixing distribution. The key idea behind the EDDP is to define a set of random measures that are marginally EDP-distributed for every possible combination of covariates $\mathbf{x}$ and $\mathbf{v}$. Using a square-breaking construction of the EDP (Wade2011), the density associated with the EDDP is
$f\quantity(u,\overline{y}\mid P)=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}w_{k}^{\phi}w_{l\mid k}^{\theta}K\quantity(u\mid\overline{y},\boldsymbol{\phi}_{k}(\mathbf{x}))K\quantity(\overline{y}\mid\boldsymbol{\theta}_{l\mid k}(\mathbf{x},\mathbf{v}))$ (6)
@@
where $w_{k}^{\phi}=v_{k}^{\phi}\prod_{l<k}(1-v_{l}^{\phi})$, $v_{k}^{\phi}\sim\mathrm{Beta}(1,\alpha_{\phi})$ with $w_{1}^{\phi}=v_{1}^{\phi}$, $w_{l\mid k}^{\theta}=v_{l\mid k}^{\theta}\prod_{j<l}(1-v_{j\mid k}^{\theta})$, $v_{l\mid k}^{\theta}\sim\mathrm{Beta}(1,\alpha_{\theta}(\phi))$ with $w_{1\mid k}^{\theta}=v_{1\mid k}^{\theta}$ for each $k$, and $\boldsymbol{\phi}_{k}(\mathbf{x})\sim P_{0\phi}^{\mathbf{x}}$, $\boldsymbol{\theta}_{l}(\mathbf{x},\mathbf{v})\sim P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}$. $\boldsymbol{\phi}_{k}(\mathbf{x})$ and $\boldsymbol{\theta}_{l\mid k}(\mathbf{x},\mathbf{v})$ are stochastic processes drawn independently from marginal distributions $P_{\phi}^{\mathbf{x}}$ and $P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\cdot\mid\phi)$, indexed by the baseline and time-varying covariates $\mathbf{x}$ and $\mathbf{v}$. This construction corresponds to the single-weights DDP model (Maceachern1999), where the atom processes are indexed by covariates, but the weights are independent of the covariates.
@@ 3.1.2 Model specifications
$\begin{split}U_{i}\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x}&\sim K_{u}(\boldsymbol{\phi}_{i}(\mathbf{x},z))\equiv\mathrm{N}\quantity((\mathbf{x},z)\boldsymbol{\beta}_{u,i}+\gamma_{i}^{z},\tau_{i}^{2}),\\ Y_{ij}\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x},\mathbf{V}_{ij}=\mathbf{v}&\sim K_{y}(\boldsymbol{\theta}_{ij}(\mathbf{x},\mathbf{v},z))\equiv\mathrm{N}\quantity((\mathbf{x},\mathbf{v},z)\boldsymbol{\beta}_{y,ij}+\psi_{ij}\gamma_{i}^{z},\sigma_{ij}^{2}),\\ \boldsymbol{\phi}_{i}(\mathbf{x},z)&\sim P_{\phi}^{\mathbf{x}},\penalty 10000\ \penalty 10000\ P_{\phi}^{\mathbf{x}}\sim\text{DP}(\alpha_{\phi},P_{0\phi}^{\mathbf{x}})\\ \boldsymbol{\theta}_{ij}(\mathbf{x},\mathbf{v},z)&\sim P_{\theta\mid\phi}^{\mathbf{x},\mathbf{v}},\penalty 10000\ \penalty 10000\ P_{\theta\mid\phi}^{\mathbf{x},\mathbf{v}}\sim\text{DP}(\alpha_{\theta\mid\phi}(\phi),P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\cdot\mid\phi)),\end{split}$ (7)
where the atom processes are expressed by simple linear models and variance parameters: $\boldsymbol{\phi}_{i}(\mathbf{x},z)=\quantity((\mathbf{x},z)\boldsymbol{\beta}_{u,i},\boldsymbol{\gamma}_{i},\tau_{i})$ and $\boldsymbol{\theta}_{ij}(\mathbf{x},\mathbf{v},z)=\quantity((\mathbf{x},\mathbf{v},z)\boldsymbol{\beta}_{y,ij},\psi_{ij},\sigma_{ij},\boldsymbol{\gamma}_{i})$. Additionally, we assume $\alpha_{\phi}\sim\mathrm{Ga}(a_{\alpha_{\phi}},b_{\alpha_{\phi}})$, $\alpha_{\theta\mid\phi}(\phi)=\alpha_{\theta\mid\phi}\sim\mathrm{Ga}(a_{\alpha_{\theta\mid\phi}},b_{\alpha_{\theta\mid\phi}})$, and the marginal base measures are $P_{0\phi}^{\mathbf{x}}=\mathrm{MVN}(\boldsymbol{\mu}_{\boldsymbol{\beta}_{u}},\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{u}})\mathrm{MVN}(\boldsymbol{\mu}_{\boldsymbol{\gamma}},\boldsymbol{\Sigma}_{\boldsymbol{\gamma}})\mathrm{IG}(a_{\tau},b_{\tau})$, where $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left(\begin{smallmatrix}\sigma_{\gamma_{0}}^{2}&\rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}\\ \rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}&\sigma_{\gamma_{1}}^{2}\end{smallmatrix}\right)$ with a given correlation parameter $\rho$, and $P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\cdot\mid\phi)=\mathrm{MVN}(\boldsymbol{\mu}_{\boldsymbol{\beta}_{y}},\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{y}})\mathrm{N}(\mu_{\psi},\sigma^{2}_{\psi})\mathrm{IG}(a_{\sigma},b_{\sigma})\delta_{\boldsymbol{\gamma}}$, with $\delta_{\boldsymbol{\gamma}}$ being a dirac measure centored on $\boldsymbol{\gamma}=(\boldsymbol{\gamma}_{1},\ldots,\boldsymbol{\gamma}_{n})$. We will discuss specific choices of the hyperparameters of each prior in the simulation studies and empirical analyses.
@@
Another salient feature of the proposed EDDPM for recurrent–terminal event data is its treatment of the subject‑specific frailty vector $\boldsymbol{\gamma}_{i}=(\gamma_{i}^{0},\gamma_{i}^{1})$. We first sample $\boldsymbol{\gamma}_{i}$ from the first‑level marginal base measure $P_{0\phi}^{\mathbf{x}}$, which governs the terminal‐event distribution, and then propagate this same value through the second‑level conditional base measure $P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\,\cdot\mid\phi)\propto\delta_{\boldsymbol{\gamma}_{i}},$ so that the nested EDDP prior carries identical frailty information into the recurrent gap‑time model. The frailty $\boldsymbol{\gamma}_{i}$ captures unobserved heterogeneity for the terminal event, while the parameters $\psi_{ij}$ determine how this heterogeneity modifies each recurrent gap time for subject $i$. This nested construction avoids the unfavorable random‐partition behavior highlighted in Section 3.1 (see also Section 3.1.3). Conditional on $\boldsymbol{\gamma}_{i}$ we assume independence between the terminal event and the recurrent events, a standard simplifying device in descriptive analyses of such data (Paulon2020; Xu2021; Tian2024). Additionally, to express the potential cross-world dependence, we take the base measure of $\boldsymbol{\gamma}_{i}$ to be $\operatorname{MVN}\!\bigl(\boldsymbol{\mu}_{\boldsymbol{\gamma}},\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\bigr)$, where $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left(\begin{smallmatrix}\sigma_{\gamma_{0}}^{2}&\rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}\\ \rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}&\sigma_{\gamma_{1}}^{2}\end{smallmatrix}\right)$. The non‑identifiable correlation between the cross‑world frailties $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$ is captured by a sensitivity parameter $\rho$, which induces the dependence structure between $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$, and hence the observed and missing potential outcomes. Although we fix $\rho$ when fitting the model, our empirical study demonstrates how varying this parameter facilitates a principled sensitivity analysis. Overall, this formulation allows for a flexible, nonparametric characterization of the survival event distribution, while still permitting subject-level random effects $\boldsymbol{\gamma}_{i}$ that may be correlated with recurrent gap times through $\psi_{ij}\boldsymbol{\gamma}_{i}$.

### v3

@@ 3.2 Identification via frailty and sensitivity analysis
Under Assumptions 14, (1) is nonparametrically identified up to distribution of the frailty $\boldsymbol{\gamma}$ as follows:
$\begin{split}\mu^{z}\quantity(t;r)&=\frac{\int_{\mathcal{X}}\int_{\Gamma}\kappa_{t,r}(z,\mathbf{x},\gamma^{z})\eta_{r}(1,\mathbf{x},\gamma^{1})\eta_{r}(0,\mathbf{x},\gamma^{0})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{\mathbf{X}}(\mathbf{x})d\boldsymbol{\gamma}d\mathbf{x}}{\int_{\mathcal{X}}\int_{\Gamma}\eta_{r}(1,\mathbf{x},\gamma^{1})\eta_{r}(0,\mathbf{x},\gamma^{0})f_{\boldsymbol{\gamma}}(\boldsymbol{\gamma})f_{\mathbf{X}}(\mathbf{x})d\boldsymbol{\gamma}d\mathbf{x}},\end{split}$ (3)
where $\Gamma$ is the support of $\boldsymbol{\gamma}$, $\kappa_{t,r}(z,\mathbf{x},\gamma^{z})=\mathbb{E}[N_{i}(t)\mid Z_{i}=z,D_{i}>r,\mathbf{X}_{i}=\mathbf{x},\gamma^{z}]$ is the rate of event occurrence among the observed survivors within each treatment group, and $\eta_{r}(z,\mathbf{x},\gamma^{z})=\mathrm{Pr}(D_{i}>r\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x},\gamma^{z})$ is the conditional survival function within each treatment group.
The proof is provided in Supplementary Material Section B. The models $\eta_{r}(z,\mathbf{x},\boldsymbol{\gamma})$ and $\kappa_{t,r}(z,\mathbf{x},\boldsymbol{\gamma})$ are induced by the models for $U_{i}$ and $\mathbf{Y}_{i}$, respctively. We detail the specification of the working models in Section 3.3. To our knowledge, this identification result for the recurrent event estimand is new and has not been established in the existing literature, including the related work of Lyu et al. (2023). Theorem 1 motivates a new g-computation strategy for estimating the target estimand based on observed data. The distribution of the frailty $f_{\boldsymbol{\gamma}}$ includes the cross-world correlation parameter $\rho$, which cannot be learned from the observed data. Because $\rho$ is non-identifiable, the effect is only partially identified. Accordingly, we treat $\rho$ as a sensitivity parameter and examine how posterior estimates of the causal estimands change across a range of plausible values, illustrated in Section 5.
Remark 1 (Frailty vs. copula for cross-world dependence in the recurrent-event setting).
Another potential strategy for cross-world modeling is the copula-based approach of Xu et al. (2022). In their time-to-event setting, the relevant within-world dependence is between two single event times and can be modeled using a bivariate kernel. By contrast, in our recurrent-event setting, the relevant within-world dependence is between an entire recurrent-event process and a terminal event, making the modeling task fundamentally more difficult because one must characterize the association between a stochastic event-history process and a survival outcome, rather than between two scalar random variables. This added complexity motivates our use of a frailty-based formulation, which offers a natural and tractable way to accommodate such dependence (see Section 3.3.2 for details).
3.3 BNP for recurrent events truncated by a terminal event
@@ 3.3.1 Enriched dependent Dirichlet process
A popular BNP approach for recurrent processes with a terminal event is jointly modeling the gap times of recurrent and terminal events using a Dirichlet process (DP) mixture model. While this is a simple approach adopted in the literature for the joint associational analysis of recurrent events with a terminal event (Paulon et al., 2020; Xu et al., 2022; Tian et al., 2024), a potential limitation is that the latent random partition induced by the DP will be overwhelmingly determined by the recurrent events rather than the terminal event as the number of recurrences grows, leading to unreliable within-cluster predictions of the survival time, with higher posterior variance due to unnecessarily small sample sizes. Additionally, existing approaches use the joint DP primarily to model a frailty term, but it is somewhat counterintuitive that the random partition, which yields the subject-specific frailty, is driven largely by numerous recurrent events rather than the subject-specific survival event. More technical discussions on this partitioning issue are provided in Supplementary Material A.
@@
We address these issues by introducing a nested structure into the DP prior tailored for the analysis of recurrent events with semi-competing risks. The key idea is to introduce a nested partition for the unknown random joint probability measure $\mathbf{P}$. In particular, we consider the random marginal $\mathbf{P}_{\phi}$ and the random conditional $\mathbf{P}_{\theta\mid\phi}$ to obtain the desired clustering structure. Then, the nested prior is defined as $\mathbf{P}_{\phi}\sim\mathrm{DP}(\alpha_{\phi},P_{0\phi})$, $\mathbf{P}_{\theta\mid\phi}\sim\mathrm{DP}(\alpha_{\theta}(\phi)$, $P_{0\theta\mid\phi}(\cdot\mid\phi))$ for all $\phi\in\Phi$, where $\mathbf{P}_{\theta\mid\phi}$ for $\phi\in\Phi$ are independent of $\mathbf{P}_{\phi}$. These assumptions induce a prior for the random joint distribution $\mathbf{P}$ through the joint law of the marginal and conditionals and the mapping $(P_{\phi},P_{\theta\mid\phi})\to\int P_{\theta\mid\phi}(\cdot\mid\phi)dP_{\phi}$. Then, the prior is parameterized by the base measure $P_{0}$, defined by $P_{0}(A\times B)=\int_{A}P_{\theta\mid\phi}(B\mid\phi)dP_{\phi}$. This nested prior is called the Enriched Dirichlet process (EDP) prior (Wade et al., 2011).
@@
One core assumption of the EDP is that the distributions are exchangeable at both survival and recurrent event levels. However, in practice, we often have access to subject-specific covariates $\mathbf{X}_{i}$, which could potentially include time-indexed covariates $\mathbf{V}_{ij}$ at the $j$-th event (e.g., event number). These variables provide useful information in characterizing the distributions at both levels, and we hence further incorporate the dependence of the covariates on the EDP through the use of the dependent Dirichlet process (DDP) (Maceachern, 1999) to relax the exchangeability assumption. We refer to the resulting prior as the enriched dependent Dirichlet process (EDDP). The enriched dependent Dirichlet process mixture (EDDPM) uses the EDDP as a prior for the mixing distribution. The key idea behind the EDDP is to define a set of random measures that are marginally EDP-distributed for every possible combination of covariates $\mathbf{x}$ and $\mathbf{v}$. Using a square-breaking construction of the EDP (Wade et al., 2011), the density associated with the EDDP is
$f\quantity(u,\overline{y}\mid P)=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}w_{k}^{\phi}(\mathbf{x})w_{l\mid k}^{\theta}(\mathbf{x},\mathbf{v})K_{u}\quantity(u\mid\overline{y},\boldsymbol{\phi}_{k}(\mathbf{x}))K_{y}\quantity(\overline{y}\mid\boldsymbol{\theta}_{l\mid k}(\mathbf{x},\mathbf{v})),$ (4)
@@
with some appropriate kernels $K_{u}$ and $K_{y}$. We introduce the covariate dependence only through atoms by letting $w_{k}^{\phi}(\mathbf{x})=w_{k}^{\phi}$ and $w_{l\mid k}^{\theta}(\mathbf{x},\mathbf{v})=w_{l\mid k}^{\theta}$, where $w_{k}^{\phi}=v_{k}^{\phi}\prod_{l<k}(1-v_{l}^{\phi})$, $v_{k}^{\phi}\sim\mathrm{Beta}(1,\alpha_{\phi})$ with $w_{1}^{\phi}=v_{1}^{\phi}$, $w_{l\mid k}^{\theta}=v_{l\mid k}^{\theta}\prod_{j<l}(1-v_{j\mid k}^{\theta})$, $v_{l\mid k}^{\theta}\sim\mathrm{Beta}(1,\alpha_{\theta}(\phi))$ with $w_{1\mid k}^{\theta}=v_{1\mid k}^{\theta}$ for each $k$, and $\boldsymbol{\phi}_{k}(\mathbf{x})\sim P_{0\phi}^{\mathbf{x}}$, $\boldsymbol{\theta}_{l}(\mathbf{x},\mathbf{v})\sim P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}$. $\boldsymbol{\phi}_{k}(\mathbf{x})$ and $\boldsymbol{\theta}_{l\mid k}(\mathbf{x},\mathbf{v})$ are stochastic processes drawn independently from marginal distributions $P_{\phi}^{\mathbf{x}}$ and $P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\cdot\mid\phi)$, indexed by the covariates $\mathbf{x}$ and $\mathbf{v}$. This construction corresponds to the single-weights DDP model (Maceachern, 1999; De Iorio et al., 2009), where the atom processes are indexed by covariates, but the weights are independent of the covariates.
@@ 3.3.2 Model specifications
$\begin{split}U_{i}\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x}&\sim K_{u}(\boldsymbol{\phi}_{i}(\mathbf{x},z))\equiv\mathrm{N}\quantity((\mathbf{x}^{\top},z)\boldsymbol{\beta}_{u,i}+\gamma_{i}^{z},\tau_{i}^{2}),\\ Y_{ij}\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x},\mathbf{V}_{ij}=\mathbf{v}&\sim K_{y}(\boldsymbol{\theta}_{ij}(\mathbf{x},\mathbf{v},z))\equiv\mathrm{N}\quantity((\mathbf{x}^{\top},\mathbf{v}^{\top},z)\boldsymbol{\beta}_{y,ij}+\psi_{ij}\gamma_{i}^{z},\sigma_{ij}^{2}),\\ \boldsymbol{\phi}_{i}(\mathbf{x},z)&\sim P_{\phi}^{\mathbf{x}},\penalty 10000\ \penalty 10000\ P_{\phi}^{\mathbf{x}}\sim\text{DP}(\alpha_{\phi},P_{0\phi}^{\mathbf{x}})\\ \boldsymbol{\theta}_{ij}(\mathbf{x},\mathbf{v},z)&\sim P_{\theta\mid\phi}^{\mathbf{x},\mathbf{v}},\penalty 10000\ \penalty 10000\ P_{\theta\mid\phi}^{\mathbf{x},\mathbf{v}}\sim\text{DP}(\alpha_{\theta\mid\phi}(\phi),P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\cdot\mid\phi)),\end{split}$ (5)
where the atom processes are expressed by simple linear models and variance parameters: $\boldsymbol{\phi}_{i}(\mathbf{x},z)=\quantity((\mathbf{x}^{\top},z)\boldsymbol{\beta}_{u,i},\gamma_{i}^{z},\tau_{i})$ and $\boldsymbol{\theta}_{ij}(\mathbf{x},\mathbf{v},z)=\quantity((\mathbf{x}^{\top},\mathbf{v}^{\top},z)\boldsymbol{\beta}_{y,ij},\psi_{ij},\sigma_{ij},\gamma_{i}^{z})$. Although we write unit- and gap-indexed latent parameters $(\boldsymbol{\phi}_{i},\boldsymbol{\theta}_{ij})$, under the enriched Dirichlet process prior, these parameters take on only finitely many distinct values almost surely. Thus, the model induces a data-adaptive clustering in which many units (and many gaps) share the same parameter values; the effective number of free parameters is the number of occupied clusters rather than $n$. Additionally, we assume $\alpha_{\phi}\sim\mathrm{Ga}(a_{\alpha_{\phi}},b_{\alpha_{\phi}})$, $\alpha_{\theta\mid\phi}(\phi)=\alpha_{\theta\mid\phi}\sim\mathrm{Ga}(a_{\alpha_{\theta\mid\phi}},b_{\alpha_{\theta\mid\phi}})$, and the marginal base measures are $P_{0\phi}^{\mathbf{x}}=\mathrm{MVN}(\boldsymbol{\mu}_{\boldsymbol{\beta}_{u}},\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{u}})\times\mathrm{MVN}(\boldsymbol{\mu}_{\boldsymbol{\gamma}},\boldsymbol{\Sigma}_{\boldsymbol{\gamma}})\times\mathrm{IG}(a_{\tau},b_{\tau})$, where $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left(\begin{smallmatrix}\sigma_{\gamma_{0}}^{2}&\rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}\\ \rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}&\sigma_{\gamma_{1}}^{2}\end{smallmatrix}\right)$ with a given correlation parameter $\rho$, and $P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\cdot\mid\phi)=\mathrm{MVN}(\boldsymbol{\mu}_{\boldsymbol{\beta}_{y}},\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{y}})\times\mathrm{N}(\mu_{\psi},\sigma^{2}_{\psi})\times\mathrm{IG}(a_{\sigma},b_{\sigma})\times\delta_{\boldsymbol{\gamma}}$, where $\delta_{\boldsymbol{\gamma}}$ denotes a Dirac measure at the frailty component $\boldsymbol{\gamma}$ contained in $\phi$, so that the same subject-level frailty is propagated from the terminal-time kernel into the recurrent-gap kernel within each top-level $\phi$-cluster. We will discuss specific choices of the hyperparameters of each prior in the simulation studies and empirical analyses.
@@
Another salient feature of the proposed EDDPM for recurrent-terminal event data is its treatment of the subject-specific frailty vector $\boldsymbol{\gamma}_{i}=(\gamma_{i}^{0},\gamma_{i}^{1})$. We first sample $\boldsymbol{\gamma}_{i}$ from the first-level marginal base measure $P_{0\phi}^{\mathbf{x}}$, which governs the terminal-event distribution, and then propagate this same value through the second-level conditional base measure $P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\,\cdot\mid\phi)\propto\delta_{\boldsymbol{\gamma}_{i}},$ so that the nested EDDP prior carries identical frailty information into the recurrent gap-time model. The frailty $\boldsymbol{\gamma}_{i}$ captures unobserved heterogeneity for the terminal event, while the parameters $\psi_{ij}$ determine how this heterogeneity modifies each recurrent gap time for subject $i$. This nested construction avoids the unfavorable random-partition behavior highlighted in Section 3.3 (see also Section 3.3.3). Conditional on $\boldsymbol{\gamma}_{i}$ we assume independence between the terminal event and the recurrent events, a standard simplifying device in descriptive analyses of such data (Paulon et al., 2020; Xu et al., 2021; Tian et al., 2024).
We parameterize cross-world dependence through the frailty-level correlation $\rho=\operatorname{Cor}(\gamma_{i}^{0},\gamma_{i}^{1})$ by placing a bivariate distribution with correlation $\rho$ in the base measure for the frailty component of the top-level EDDP atom. Concretely, the $\boldsymbol{\gamma}$-margin of the top-level base measure is $\operatorname{MVN}\!\bigl(\boldsymbol{\mu}_{\boldsymbol{\gamma}},\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}\bigr)$, where $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}=\left(\begin{smallmatrix}\sigma_{\gamma_{0}}^{2}&\rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}\\ \rho\sigma_{\gamma_{0}}\sigma_{\gamma_{1}}&\sigma_{\gamma_{1}}^{2}\end{smallmatrix}\right)$. Because the EDDP is centered on its base measure, the induced prior predictive marginal law of a randomly selected subject’s frailty vector $\boldsymbol{\gamma}_{i}$ equals this $\boldsymbol{\gamma}$-margin, and therefore $\operatorname{Cor}(\gamma_{i}^{0},\gamma_{i}^{1})=\rho$ under the prior. In our enriched construction, $\boldsymbol{\gamma}_{i}$ is sampled at the top level (through the $\phi$-atoms governing the terminal event) and then propagated to the recurrent-event kernel via the conditional base measure $P_{0\theta\mid\phi}^{\mathbf{x},\mathbf{v}}(\cdot\mid\phi)\propto\delta_{\boldsymbol{\gamma}_{i}}$, so that the same subject-level frailty enters both the terminal- and recurrent-event components within each top-level cluster.
Inference is performed conditional on a fixed $\rho$, which is not identified from the observed data because $(\gamma_{i}^{0},\gamma_{i}^{1})$ are never jointly observed. Accordingly, we treat $\rho$ as a sensitivity parameter and examine the robustness of posterior inferences for the causal estimands over a plausible grid of $\rho$ values. We emphasize that $\rho$ governs cross-world association at the frailty level, inducing the dependence structure between potential outcomes, but the induced association (e.g., $U_{i}^{0}$ and $U_{i}^{1}$) may differ due to additional outcome-specific variability and model components beyond $\boldsymbol{\gamma}_{i}$. Overall, this formulation allows for a flexible, nonparametric characterization of the survival event distribution, while still permitting subject-level random effects $\boldsymbol{\gamma}_{i}$ that may be correlated with recurrent gap times through $\psi_{ij}\boldsymbol{\gamma}_{i}$.